Accelerator Update for May 2, 2008 - May 5, 2008
Friday May 2
The day shift began with Operations monitoring the Tevatron (TeV) store 6109 (120.24E30) and the Recycler stash (60.2E10), with the Antiproton Source stacking (34.1E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with NuMI, MTest and the SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 10:57 AM, the first of several power glitches occurred. Due to various problems caused by the glitches, 30 minutes of beam time was lost.
Operations terminated store 6109 at 4:03 PM.
Beam was turned off to MTest at 6:05 PM.
At 7:30 PM, the Pbar permit dropped due to control problems with the water skid for the Pbar lens. Experts were consulted.
Pbar controls problems were corrected by 9:30 PM.
HEP shot setup began at 9:35 PM.
Shot setup was put on hold at 10:17 PM, the C0 Cooling pump tripped off thus causing the 4 Cryo compressors at C0 to trip off. The Duty Mechanic and Cryo experts are investigating.
Saturday May 3
The midnight shift began with shot setup in standby, with the stash at (270.1 E10), with Pbar stacking (34.0E10), and with NuMI and BNB taking beam.
At 1:49 AM, the C0 pump was back on (with reduced pressure), and one Cyro compressor (C0 #4) has been on for the past 45 minutes. We normally run with four compressors at C0, but things appear to be stable enough to go into shot setup.
HEP shot setup began at 2:48 AM.
Operations established store 6113 at 4:26 AM with an initial luminosity of 211.73E30.
MTest turned on at 6:31 AM.
C0 #4 compressor tripped off at 7:37 AM, since Cryo temperatures at C0 seemed pretty stable it was decided to continue to run like this.
Operations terminated store 6113 at 3:05 PM.
Operations prepared D0 for an access at 3:19 PM. They came out of access at 4:38 PM.
At 4:51 PM, a fire hose was run from the C0 assembly building to provide cooling water at C0. Cryo was then able to turn on the C0 #4 compressor.
Beam was turned off to MTest at 5:28 PM.
HEP shot setup began at 5:37 PM.
Operations established store 6115 at 7:20 PM with an initial luminosity of 255.58E30.
At 8:05 PM, a high energy Linac quad bulk power supply tripped off, ops were unable to reset it. An Expert was called and is on his way in.
The expert found and replaced the bad power supply cards, beam returns at 9:58 PM.
Sunday May 4
The midnight shift began with Operations monitoring store 6115 and the stash, with Pbar stacking, and with NuMI and BNB taking beam.
At 2:23 AM a power glitch aborted the Tev beam.
HEP shot setup began at 4:20 AM.
Operations established store 6117 at 5:57 AM with an initial luminosity of 194.61E30.
MTest turned on at 6:52 AM.
Beam was turned off to MTest at 6:28 PM.
Operations terminated store 6117 at 9:50 PM.
HEP shot setup began at 10:03 PM.
Operations established store 6118 at 11:52 PM with an initial luminosity of 264.37E30.
Monday May 5
The midnight shift began with Operations monitoring store 6118 and the stash, with Pbar stacking, and with NuMI and BNB taking beam.
MTest turned on at 7:04 AM.
The Plans for Monday
To increase the size of the hose that is feeding the C0 compressors from 2.5 in. diameter to 4.0 in. diameter. Continue to stack, stash and store.
Comments and Suggestions
|last modified 05/05/2008 email Fermilab|
|Security, Privacy, Legal|